Time Warner shares have a market capitalization of $50 billion. The company is expected to pay a dividend of $0.50 per share and each share trades for $30. The growth rate in dividends is expected to be 5% per year. Also, Time Warner has $20 billion of debt that trades with a yield to maturity of 8%. If the firm's tax rate is 40%, what is the WACC?
Answer #1

Cost of equity = (Dividend / Price) + Growth rate = ($0.50 / $30) + 0.05 = 0.0667 or 6.67%

Cost of debt = Yield to maturity(1 - Tax rate) = 0.08(1 - 0.40) = 0.048 or 4.8%

WACC = [0.0667 × ($50 billion / $70 billion] + [0.048 × ($20 billion / $70 billion] = 0.0613 or 6.13%